### v1

# Additional dimensions of space and time in the region of deep inelastic processes

Abstract
We prove that the well-known Heisenberg uncertainty relations and Landau-Peierls uncertainty relations implicitly contain “hidden” angular variables, which belong to new uncertainty relations. Based on the obtained relations, we derive a formula for estimating the speed $U^{*}$ of a virtual particle in indirect measurements. We applied the theory of indirect measurements and the derived formula to estimate the module of the group velocity of virtual photons from the DIS HERA data. The HERA data indicate that the speed of virtual photons exceeds the speed of light $c$ in free space, $U^{*}>c$. The properties of virtual photons and a hypothetical tachyon particle are almost identical. It is found that in the realm of particle interaction, the new angular parameters are closely related to the type of the phase-space geometry and dimensionality of the space-time continuum. It is suggested that the problem of the normalization condition $U^{*}=c$ at $Q^{2}=0\,\rm{GeV}^{2}$ can be solved naturally within the framework of “Two-Time Physics” developed by I. Bars. 2T-physics is the theory with local symplectic $\mathrm{Sp(2,R)}$ gauge symmetry in phase-space and the space-time geometry of signature $\mathrm{(1+1^{\prime},d+1^{\prime})}$ with one extra time-like and one extra space-like dimensions.
I Introduction
Physical processes in quantum systems can only be properly described by introducing various types of intermediate states. Scattering processes of real particles are also described by the quantum mechanism of exchange of virtual particles (gauge bosons, resonance states, more complex objects, such as Regge trajectories). The concept of a virtual off-mass-shell particle is derived from the microscopic violation of causality allowed by the time-energy uncertainty relation [1, 2, 3, 4, 5]. The relation between the momentum and energy of a virtual particle can be anything that is required by the conservation of 4-momentum at the vertices. It should be noted that the content of the term “virtual particle” has undergone a significant change. Even in the recent past, virtual particles usually meant such particles in virtual states (e.g. photons, electrons, pions) that were well studied in real states. A class of particles (quarks, gluons, etc.) has emerged which, due to the confinement property of quantum chromodynamics, cannot in principle be in real states.
Although it is impossible to observe such intermediate states directly, their experimental study is of great interest and importance due to their nontrivial dynamical properties. Nevertheless, a number of properties of the virtual particle can be measured indirectly. Indirect measurement is a measurement in which the value of the unknown quantity sought is calculated from measurements of other quantities related to the measurand by some known relation [6, 7].
The goal of the present report is twofold. First, we derive a formula for estimating a speed of virtual particles. Second, we combine the theory of indirect measurements, the hardware resolution of the ZEUS detector [8, 9] and the obtained formula into a mathematical tool for evaluating the speed of virtual particles from experimental data. As an application, we present preliminary results on the speed of virtual photons ($\gamma^{*}$) using a small set of deep inelastic scattering (DIS) data from the HERA collider [10].
Natural questions arises: What is the value of data on the speed of virtual particles? What new information would they reveal? Especially considering that at relativistic velocities, a particle’s speed is less informative than its momentum. And in the case of a real photon, knowing only its speed tells us nothing about the photon’s energy or momentum.
Let us recall three phenomena: 1) The formation of a Mach cone when the speed of a body in a medium is close to or greater than the speed of sound in the medium; 2) The Cherenkov-Vavilov effect [11, 12, 13]; 3) In 1904-1905 A. Sommerfeld [14] established in the context of Lorentz’s theory of electromagnetism [15] that if the speed of the electron $\mathrm{U}$ is less than the speed of light in a vacuum, $\mathrm{U}<c$, then the electron is able to move at a constant velocity. However, if $\mathrm{U}>c$, then an external force is required for uniform motion.
What these examples have in common is that when the particle speed exceeds a certain characteristic value, $\mathrm{U}_{cr}$, the dynamics of the process changes dramatically. Moreover, the presence of particles with velocities above the critical value serves as a natural indicator of changes in the properties of the phase space itself in which the process is taking place.
The “implementation” of the uncertainty principle in nature determines the existence of virtual particles with a very wide range of dynamic properties. For this reason, the present derivation of a formula for the lower bound on the speed of a virtual particle is based on the Heisenberg’s uncertainty relations (HUR).
The uncertainty principle and the uncertainty relations for observables of two canonically conjugate quantum mechanical operators discovered by Heisenberg [1] are fundamental foundations of quantum mechanics. In the article [1], only a heuristic estimate was given of how the inaccuracy of the particle coordinate, $q_{1}$, is associated with the inaccuracy of the particle momentum, $p_{1}$, into one relation, $p_{1}q_{1}\sim\hbar$, called the uncertainty relation ( see also Refs [16], [17], [18]). H. Weyl [19] provide another proof of HUR and also gave the inequality a modern look,
$\Delta p_{x}\Delta x\geq\hbar/2.$ (1)
In the next two paragraphs we present derivations of new uncertainty relations and a formula for evaluating the speed of a virtual particle in indirect measurements. After presenting experimental results for the lower bound on the speed of virtual photons, we discuss the tachyonic properties of $\gamma^{*}$, followed by arguments that allow relating the superluminal speed of virtual photons to the metric properties of space-time (in particular, its dimensionality).
II New angular variables and uncertainty relations
Uncertainties of quantum mechanical Hermitian operators $\hat{x}$ and $\hat{p_{x}}$ are defined (see Ref. [19], p. 77 and Ref. [20], p. 137) via
[…]
Therefore, we write out the uncertainty relations for all projections of the pair of conjugate coordinate-momentum pair in terms of mean square deviations:
$(\Delta p_{x})^{2}(\Delta x)^{2}$$\geq$ $(\hbar/2)^{2},$
$(\Delta p_{y})^{2}(\Delta y)^{2}$$\geq$ $(\hbar/2)^{2},$ (3)
$(\Delta p_{z})^{2}(\Delta z)^{2}$$\geq$ $(\hbar/2)^{2},$
without extracting the square root, as in (1). If we now add the left-hand sides of these inequalities, we find that the resulting sum is a dot product $\mathbf{(\Delta P)^{(2)}\cdot(\Delta R)^{(2)}}$ of the vectors $\mathbf{(\Delta P)^{(2)}}=((\Delta p_{x})^{2},(\Delta p_{y})^{2},(\Delta p_{z})^{2})$ and $\mathbf{(\Delta R)^{(2)}}=((\Delta x)^{2},(\Delta y)^{2},(\Delta z)^{2})$. The dot product specifies the angle between vectors, and therefore for the norm of vectors $\mathbf{(\Delta P)^{(2)}}$ and $\mathbf{(\Delta R)^{(2)}}$ we obtain the uncertainty relation, which includes a new angular variable $\psi$:
$\|\mathbf{(\Delta P)^{(2)}}\|\,\|\mathbf{(\Delta R)^{(2)}}\|\geq\frac{3\hbar^{2}}{4\cos\psi}.$ (4)
The relations (3) consist of only positive definite terms and this defines the domain of the angle $\psi\,\in[0,\pi/2)$, and the domain of the function values, $0<\cos{\psi}\leq 1$. Thus, depending on the state of the physical system under study, the value of $\cos{\psi}$ varies and imposes constraints on $\|\mathbf{(\Delta P)^{(2)}}\|$ and $\|\mathbf{(\Delta R)^{(2)}}\|$ of different degrees of stiffness. The function $\cos{\psi}$ appears as a result of reducing the six degrees of freedom in (3) to three degrees of freedom in (4).
III Estimation of a particle speed from the uncertainty relations
In the same 1927 article [1], Heisenberg gives an uncertainty relation for another pair of canonically conjugate energy-time variables. This relation is only definite up to Planck’s constant, so we write it out by including an arbitrary constant $\delta_{H}$:
$(\Delta E)^{2}(\Delta t)^{2}\geq{\delta^{2}\!\!\!}_{{}_{H}}\,\hbar^{2},$ (5)
whose value is fixed by the conditions of the problem being solved. Landau and Peierls [2] generalized a number of conclusions from classical quantum mechanics to the relativistic domain. In particular, it was shown that the Heisenberg inequalities for momentum and coordinates are also valid at relativistic velocities. In passing to the relativistic consideration, however, the inequality (5) does not give such a simple justification. Nevertheless, Landau and Peierls have derived new inequalities for a free relativistic particle, Refs. [2, 21]:
$|\mathrm{U}_{i}|\Delta p_{i}\Delta t\geq{\delta\!}_{{}_{LP}}\hbar,$ (6)
that holds for each of the components $i=(x,y,z)$ separately. Here the symbol $\mathbf{U}$ denotes the group velocity vector of the particle, $\mathbf{U}=(\mathrm{U}_{x},\mathrm{U}_{y},\mathrm{U}_{z})$ and an arbitrary constant $\delta_{{}_{LP}}$ is introduced on the same reasoning as in the inequality (5). Adding the squares of the relations (6) for $i=(x,y,z)$, as above, we get on the left-hand side of the inequality the scalar product $\mathbf{U^{(2)}\cdot(\Delta P)^{(2)}}$ of vectors $\mathbf{U^{(2)}}=((\mathrm{U}_{x})^{2},(\mathrm{U}_{y})^{2},(\mathrm{U}_{z})^{2})$ and $\mathbf{(\Delta P)^{(2)}}$. Thus we reveal another “hidden” angle ${\psi}_{{}_{H}}$ between the phase-space vectors and obtain another inequality connecting the norms of the particle’s quadratic velocity vector, the mean square deviation of its momentum and the square of the duration of the measurement process:
$\|\mathbf{U^{(2)}}\|\,\|\mathbf{(\Delta P)^{(2)}}\|(\Delta t)^{2}\geq 3(\delta_{{}_{LP}}\hbar)^{2}/\cos{\psi}_{{}_{H}}.$ (7)
Using inequalities (4) and (7), we are now able to estimate the module of the particle group velocity $|\mathbf{U}|$ in conditions where the direct measurement of the velocity is impossible (the method of indirect measurements [6]). For this purpose, the ratio of the inequality (7) to (4) or the ratio of the inequality (7) to (5), respectively, must be taken. In this way,
$\|\mathbf{U^{(2)}}\|\geq A_{t}\frac{(\Delta E)^{2}}{\|\mathbf{(\Delta P)^{(2)}}\|}.$ (8)
Finally, by means of the Cauchy-Buniakowsky-Schwarz inequality, we obtain the following estimate of the lower bound of the norm of the velocity,
$\|\mathrm{U}_{lb}^{*}\|\sim\sqrt{\sqrt{3}\|\mathbf{U^{(2)}}\|}=\sqrt{\sqrt{3}A_{t}\frac{(\Delta E)^{2}}{\|\mathbf{(\Delta P)^{(2)}}\|}}.$ (9)
Here $A_{t}=3{\delta^{2}\!\!\!}_{{}_{LP}}/({\delta^{2}\!\!\!}_{{}_{H}}\,\cos{\psi}_{{}_{H}})$ is the theoretical magnitude of the normalization parameter. In the next paragraph, we will see how the velocity of virtual particles relates the value of $A_{t}$ and the metric properties of space-time.
IV Space-time metrics in the interaction region
To classify (pseudo-)Euclidean spaces, the so-called space index $k$ (or the index of inertia) is introduced. It is defined as the number of imaginary unit basis vectors of the orthonormal frame [22, 23]. For the proper Euclidean space, $n=d$, the space index $k=0$. For the Minkowski space with the total dimension $n=d+1=4$ and the signature (+,-,-,-)=(1, 3), the space index of $k=3$.
Figure 1: (a) Kinematic ($x_{{}_{Bj}}$, $Q^{2}$) grid of the combined HERA I data for neutral current $e^{+}p$ deep inelastic scattering events at the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Different colors indicate data with the inelasticity, $y=Q^{2}/x_{{}_{Bj}}{s}_{com}$, in the following intervals : $0.354-0.47$ (red); $0.47-0.511$ (blue); $0.511-0.598$ (black); $0.598-0.676$ (magenta); $0.676-0.951$ (green). (b) The normalized virtual photon speed, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of the photon virtuality $Q^{2}$. The color of the dots corresponds to different values of the inelasticity $y$ as in Fig (a). It follows from the condition, $\beta^{*}(Q^{2}\rightarrow 0)=$1 that the normalization parameter $A_{e}=1/\sqrt{3}$. (c) The same as in panel (b), but for $\beta^{*}$ as a function of two variables $x_{{}_{Bj}}$ and $Q^{2}$ at different $y$-intervals. (d) The same as in panel (c), but the figure is rotated so that the data points are roughly projected onto a curve.
Let us first discuss possible values of the parameter $A_{t}$ in formula (9), and denote by $A_{e}$ its value found from the data. The Standard Model assumes that the geometry of space-time known from macroscopic physics also holds in the microcosm too. The derivation of the inequality (4) was based on this assumption, using the dot product of vectors in proper Euclidean space. The inequality (6) was derived from the inequality (5). Therefore, there is a good reason to believe that ${\delta^{2}\!\!\!}_{{}_{LP}}={\delta^{2}\!\!\!}_{{}_{H}}$. In this case, $A_{t}=3/\cos{\psi}_{{}_{H}}\geq 3$. Consequently, if it follows from an experimental data that $A_{e}\geq 3$, then the interaction of particles takes place in the region of space-time with the Minkowski geometry and the space index of $k=3$.
The case is quite different if $A_{e}<3$. Then our assumption about the metric of space-time in the interaction region is not correct.
Let us now apply the formula (9) to evaluate the speed of virtual photons. In DIS processes $c|\vec{q}|=\sqrt{q_{0}^{2}+Q^{2}}$ and therefore the quantities $(\Delta E)^{2}=(\Delta q_{0})^{2}$ and $\|\mathbf{(\Delta P)^{(2)}}\|=(\Delta q)^{2}$ depend on the kinematic variables $x_{{}_{Bj}}$, $y$, $Q^{2}$ and uncertainties of their measurements by the following chain of relations,
$(\Delta q)^{2}$ $=$ $\frac{1}{c^{4}}\Big{[}\Big{(}\frac{q_{0}}{|\vec{q}|}\Big{)}^{2}(\Delta q_{0})^{2}+\frac{(\Delta Q^{2})^{2}}{4(\vec{q})^{2}}\Big{]},$ (10)
$(\Delta q_{0})^{2}$ $=$ $c^{2}P^{2}y^{2}(\Delta x_{{}_{Bj}})^{2}$ (11)
$+\,c^{2}(l-x_{{}_{Bj}}P)^{2}(\Delta y)^{2},$
$(\Delta Q^{2})^{2}$ $=$ $\frac{4Q^{2}}{1-y}(c\Delta p_{t})^{2}+\Big{(}\frac{Q^{2}}{1-y}\Big{)}^{2}(\Delta y)^{2},$ (12)
$(\Delta x_{{}_{Bj}})^{2}$ $=$ $\frac{4x_{{}_{Bj}}^{2}}{(1-y)Q^{2}}(c\Delta p_{t})^{2}$ (13)
$+\,x_{{}_{Bj}}^{2}\Bigg{[}\frac{1}{(1-y)^{2}}+\frac{1}{y^{2}}\Bigg{]}(\Delta y)^{2}.$
The chain of these relations is closed if to enter the resolution of the central tracking detector $\sigma(p_{t})/p_{t}$ and the energy resolution of the uranium calorimeter $\sigma(\Sigma_{e})/\Sigma_{e}$ [8, 9].
As input, we use the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider [10]. The magnitude of the photon virtuality, $Q^{2}$, varies over a very wide range of values. Neutral current interaction cross sections at low $Q^{2}\leq$100 GeV2 are dominated by virtual photon exchange. In the limit $Q^{2}\rightarrow 0$ GeV2 (a real photon limit), $\beta^{*}=\|\mathrm{U}_{lb}^{*}/c\|\rightarrow 1$ should hold. This condition allows to fix the value of $A_{e}$ and at $Q^{2}>0$ GeV2 to set a lower bound on the speed of virtual photons.
The kinematic range of the combined HERA I data with $Q^{2}\leq$100 GeV2 is shown in Fig. 1(a). Data points are grouped into strips with similar values of inelasticity $y$ and marked with different colors (see figure caption). Such structuring reflects the kinematic relationship between the variables $x_{{}_{Bj}}$, $y$ and $Q^{2}$, and the procedure for combining data from two different experiments by translation onto common grids [10].
Figure 1(b) shows the speed of virtual photons normalized to the speed of real photons, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of $Q^{2}$ at different $y$-intervals. The result follows from Eqs. (9) - (13) and the HERA I data for neutral current $e^{+}p$ deep inelastic scattering events with beam momenta $(l,P)=$(27.5, 820) GeV/c and the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Note that $\beta^{*}$ grows with $Q^{2}$, but the slope of this growth decreases with $y$. The condition $\beta^{*}(Q^{2}\rightarrow 0)=$1 fixes $A_{e}$. In this way we get $\sqrt{3}A_{e}=1$. Figures 1(c) and 1(d) show the dependence of $\beta^{*}$ on two variables, $x_{{}_{Bj}}$ and $Q^{2}$. And again we see that $\beta^{*}$ grows with $Q^{2}$ and $x_{{}_{Bj}}$. Figure 1(d) is rotated view of Fig. 1(c) in order to project all data points on the same curve. This demonstrates that almost all points at different $y$ are located on a flat surface.
The results presented in Fig. 1 show that the speed of virtual photons at $Q^{2}>0$ GeV2 exceeds the speed of light $c$ in free space, $\beta^{*}>1$. Superluminal virtual photons can be interpreted as representatives of a class of superluminal particles, the tachyons [24, 25, 26, 27, 28, 29], since their properties are largely similar as shown in Table 1,
Table 1: Comparison of the properties of tachions and virtual photons.
Tachyon Virtual photon, $\gamma^{*}$
mass $(m^{*})^{2}<0$ $(m^{*})^{2}=-Q^{2}<0$
energy $\epsilon<0,>0$ $q_{0}<0,>0$
speed $\beta^{*}\geq 1$ $\beta^{*}\geq 1$
$\epsilon\rightarrow 0,\beta^{*}\rightarrow\infty$ $q_{0}\rightarrow 0,\beta^{*}\rightarrow\infty$ (?)
From the above discussion, the value of $A_{e}=1/\sqrt{3}<3$ obtained from the experimental data, contradicts the theoretical expectation, $A_{t}=3/\cos{\psi}_{{}_{H}}>3$. Therefore, we want to identify a theoretical scenario that can help eliminate this contradiction.
I) Nonlinear fields in 4D [30, 31]. In nonlinear field theories (with scalar or/and vector fields) the signal velocity depends on the magnitude of the field and its derivatives. In this case, the speed of the signals can be less or more than the speed of light $c$ in void and the space index change from $k=3$ to $k=2$ or even to $k=1$ and hence, $A_{t}<3$. This corresponds to the space-time domain, in which there are two or three time-like dimensions. However, in this scenario there are problems not only with causality (paradoxes), the violation of unitarity and occurrence of ghost modes with negative norm (for more discussion see Refs. [32, 33]), but also the entire relativistic kinematics in the Minkowski space is destroyed.
II) Stochastic time. One can increase the pseudo-Euclideanity of space-time by adding an additional time-like dimension in the interaction region, i.e., going from signature (1,3) to signature (2,3). Analysis of this scenario shows [33], that in the case of a space-time with a thermally excited second time dimension the dynamics of the physical system is diffusive, not ballistic, so that all trajectories involving motion in the second time dimension are dynamically unstable, thus allowing us to avoid the difficulties outlined for nonlinear field theories. The thermal extra time dimension behaves like an extra spatial dimension and the high temperature limit for the thermal extra time dimension is equivalent to a small compactification radius for an extra spatial dimension. Therefore, such a scenario is consistent with Kaluza-Klein type theories. However, simply adding another spatial or temporal dimension does not solve the problem we encountered ($A_{t}>3$, $A_{e}<3$).
III) Two-time physics. Now we have to account that the vectors $\mathbf{X}$, $\mathbf{P}$ and $\mathbf{(\Delta R)^{(2)}}$, $\mathbf{(\Delta P)^{(2)}}$ are elements of the phase space. In the geometric approach, the phase space $(\mathbf{Q}^{i},\mathbf{P}_{i})$ of classical mechanics as well as of quantum mechanics is taken to be a smooth manifold equipped with a symplectic form, which induce a so-called symplectic geometry [34]. Note that in the early 60’s attempts were made to systematize hadrons with the use of the symplectic group Sp(6) [35, 36], and in the 80’s gauge theories with Higgs models based on simple classical Lie groups [37] and, in particular, the symplectic group Sp(m) were investigated.
The principle of local gauge invariance is an important ingredient in the construction of realistic models for the interactions observed in nature. I. Bars with co-authors [38], [39] [32] discovered a fundamental role of the local symplectic Sp(2, R) gauge symmetry in phase space, which gave rise to new field theories in (2,4) space-time with one extra time and one extra space dimensions (2T-physics). This includes 2T field theories that yield one-time (1T) field theories for the Standard Model [40], General Relativity [41], SUSY [42] and others. The canonical transformations of the type Sp(2, R), considered as a local gauge symmetry, has the power to cure the ghost and causality problems of extra time-like dimensions.
In the following, we give a qualitative argument for the fact that the condition $A_{t}<3$ is feasible in the framework of 2T-physics. First, we postulate that the interaction region of DIS has a small spatial and temporal extent, and within the framework of the 2T-physics, the interactions take place in $\mathrm{(1+1^{\prime},3+1^{\prime})}$ space with one extra time and one extra space dimensions and the gauge invariant sector of 2T-physics, namely the ghost free physical sector, effectively becomes a 1T theory with an effective 1+3 dimensions [43]. Mathematically, the speed of a photon in the physical sector is a combination of the three effective spatial dimensions and one effective temporal dimension (the ”effective Minkowski space-time”). So, we assume that during the DIS process a virtual photon “captures” (or “admixes”) “new” additional spatial $x$ or time-like $\tau$ dimensions with the probability $\alpha$ and captures the “old” space-time dimensions, $\{t,x^{i}\}$, with the probability $\omega$. We label the components containing the conventional time dimension with index $H$, and the components containing an additional time-like dimension with index $B$. As a result, the squares of the normalized photon velocities as spatio-temporal combinations give contributions with the following probabilities: ${\beta^{2}}_{H}(t,x^{1},x^{2},x^{3})\sim\omega$, $\tilde{\beta^{2}}_{H}(t,x^{i},x^{j},x)\sim 3\alpha$, ${u^{2}}_{B}(\tau,x^{1},x^{2},x^{3})\sim\alpha$ and $\tilde{u^{2}}_{B}(\tau,x^{i},x^{j},x)\sim 3\alpha^{2}$. The total probability must satisfy the condition:
$\omega+4\alpha+3\alpha^{2}=1.$ (14)
In particular, $\alpha=0$ if $\omega=1$, in accordance with the definition of the introduced probabilities.
Let us now find out how the parameter $A$ changes when additional dimensions are taken into account, and denote it as $A_{t}=A_{eff}$. Then, in accordance with Eq. (9) one get
${\beta^{*}}^{2}_{eff}=\sqrt{3}(\omega{\beta^{2}_{H}}A_{H}+3\alpha\tilde{\beta}^{2}_{H}A_{H}$
$+\alpha{u^{2}_{B}}A_{B}+3\alpha^{2}\tilde{u}^{2}_{B}A_{B}).$ (15)
Here, $A_{H}=3/\cos{\psi}_{{}_{H}}$ and $A_{B}=3/\cos{\psi}_{{}_{B}}$.
In the limit $Q^{2}\to 0$, all speeds in Eq. (15) should be equated to unity. As a result, we get the normalization condition $\sqrt{3}A_{eff}=1$,
$1=3\sqrt{3}\Big{[}\frac{1-\alpha-3\alpha^{2}}{\cos{\psi}_{{}_{H}}}+\frac{\alpha+3\alpha^{2}}{\cos{\psi}_{{}_{B}}}\Big{]}.$ (16)
By solving this equation for $\alpha$, we find that $\alpha>0$ if $\cos{\psi}_{{}_{B}}>\cos{\psi}_{{}_{H}}$ and $\alpha<1$ if the following condition is met
$(12\sqrt{3}-\cos{\psi}_{{}_{B}})\cos{\psi}_{{}_{H}}<9\sqrt{3}\cos{\psi}_{{}_{B}}.$ (17)
These restrictions on $\cos{\psi}_{{}_{H}}$ and $\cos{\psi}_{{}_{B}}$ are quite soft. Thus, in the space-time volume of a high-energy reaction a “mixing in” of extra dimensions is possible and the normalization condition $\sqrt{3}A_{eff}=1$ is feasible. The specific values of $\cos{\psi}_{{}_{H}}$, $\cos{\psi}_{{}_{B}}$, $\alpha$, ${\beta^{2}}_{H}$, $\tilde{\beta^{2}}_{H}$, $\tilde{u^{2}}_{B}$, $\tilde{u^{2}}_{B}$ can be found by fitting Eq. (15) to the data similar to Fig. 1(c).
V Conclusions
In this Letter, we have developed the method based on the quantum uncertainty relations and the theory of indirect measurements, which allows to evaluate the module of the group velocity of virtual photons $\gamma^{*}$ and to reveal the type of space-time geometry in the interaction region of high-energy deep inelastic scattering processes. The HERA data show that the speed of virtual photons, $\beta^{*}$, exceeds the speed of light in free space, $\beta^{*}\geq 1$. The superluminal speed means that the virtual photons $\gamma^{*}$ behave tachyon-like. A solution of the problem of the normalization condition $\beta^{*}=1$ at $Q^{2}=0\,\rm{GeV}^{2}$ is proposed in the framework of ideas of ”2T-physics” developed by I. Bars. It is shown that the “admixture” of extra time and extra space dimensions in DIS processes makes it possible to satisfy the normalization condition $\sqrt{3}A_{eff}=1$ as observed in the effective physical (1,3) Minkowski space-time.
The experimental results for virtual photons presented here are based on only a small fraction of the HERA data to illustrate the method, but the analysis can be extended to the full data set, and these investigations are left to future work.
ACKNOWLEDGMENTS
The author thanks M. Wing for reading an earlier version of the manuscript and suggestions, E. Boos, L. Dudko, A. Geiser, N. Nikitin, E. Oborneva, I. Volobuev and other colleagues from SINP MSU and the ZEUS collaboration for discussions and comments. The author is especially grateful to K. G. Gulamov, N. N. Nikolaev and B.S. Yuldashev for their mentorship in the early days. The author expresses his gratitude to the Ministry of Science and Education of the Russian Federation and the DESY Directorate for their long-term financial support and for the hospitality extended to the non-DESY members of the ZEUS collaboration.

### v2

@@
-Additional dimensions of space and time in the region of deep inelastic processes
+Additional dimensions of space and time in the domain of deep inelastic processes
 Abstract
 We prove that the well-known Heisenberg uncertainty relations and Landau-Peierls uncertainty relations implicitly contain “hidden” angular variables, which belong to new uncertainty relations.
-Based on the obtained relations, we derive a formula for estimating the speed $U^{*}$ of a virtual particle in indirect measurements.
+Based on the obtained relations, we derive a formula for estimating speed $U^{*}$ of a virtual particle in indirect measurements.
 We applied the theory of indirect measurements and the derived formula to estimate the module of the group velocity of virtual photons from the DIS HERA data.
@@
 The properties of virtual photons and a hypothetical tachyon particle are almost identical.
-It is found that in the realm of particle interaction, the new angular parameters are closely related to the type of the phase-space geometry and dimensionality of the space-time continuum.
-It is suggested that the problem of the normalization condition $U^{*}=c$ at $Q^{2}=0\,\rm{GeV}^{2}$ can be solved naturally within the framework of “Two-Time Physics” developed by I.
+It is found that in the realm of particle interaction, the new angular parameters are closely related to the type of the phase-space geometry and dimensionality of the spacetime continuum.
+It is suggested that the problem of the normalisation condition $U^{*}=c$ at $Q^{2}=0\,\rm{GeV}^{2}$ can be solved naturally within the framework of “Two-Time Physics” developed by I.
 Bars.
-2T-physics is the theory with local symplectic $\mathrm{Sp(2,R)}$ gauge symmetry in phase-space and the space-time geometry of signature $\mathrm{(1+1^{\prime},d+1^{\prime})}$ with one extra time-like and one extra space-like dimensions.
+2T-physics is the theory with local symplectic $\mathrm{Sp(2,R)}$ gauge symmetry in phase-space and the spacetime geometry of signature $\mathrm{(1+1^{\prime},d+1^{\prime})}$ with one extra time-like and one extra space-like dimensions.
 I Introduction
@@
 Indirect measurement is a measurement in which the value of the unknown quantity sought is calculated from measurements of other quantities related to the measurand by some known relation [6, 7].
-The goal of the present report is twofold.
+The goal of the present communication is twofold.
 First, we derive a formula for estimating a speed of virtual particles.
@@
 However, if $\mathrm{U}>c$, then an external force is required for uniform motion.
+It is not superfluous to note that most of these conclusions were made much earlier by O.
+Heaviside [16, 17].
 What these examples have in common is that when the particle speed exceeds a certain characteristic value, $\mathrm{U}_{cr}$, the dynamics of the process changes dramatically.
@@
 $\Delta p_{x}\Delta x\geq\hbar/2.$ (1)
-In the next two paragraphs we present derivations of new uncertainty relations and a formula for evaluating the speed of a virtual particle in indirect measurements.
-After presenting experimental results for the lower bound on the speed of virtual photons, we discuss the tachyonic properties of $\gamma^{*}$, followed by arguments that allow relating the superluminal speed of virtual photons to the metric properties of space-time (in particular, its dimensionality).
+In the next two sections we present derivations of new uncertainty relations and a formula for evaluating the speed of a virtual particle in indirect measurements.
+After presenting experimental results for the lower bound on the speed of virtual photons, we discuss the tachyonic properties of $\gamma^{*}$, followed by arguments that allow relating the superluminal speed of virtual photons to the metric properties of spacetime (in particular, its dimensionality).
 II New angular variables and uncertainty relations
@@
 $\|\mathrm{U}_{lb}^{*}\|\sim\sqrt{\sqrt{3}\|\mathbf{U^{(2)}}\|}=\sqrt{\sqrt{3}A_{t}\frac{(\Delta E)^{2}}{\|\mathbf{(\Delta P)^{(2)}}\|}}.$ (9)
-Here $A_{t}=3{\delta^{2}\!\!\!}_{{}_{LP}}/({\delta^{2}\!\!\!}_{{}_{H}}\,\cos{\psi}_{{}_{H}})$ is the theoretical magnitude of the normalization parameter.
-In the next paragraph, we will see how the velocity of virtual particles relates the value of $A_{t}$ and the metric properties of space-time.
-IV Space-time metrics in the interaction region
+Here $A_{t}=3{\delta^{2}\!\!\!}_{{}_{LP}}/({\delta^{2}\!\!\!}_{{}_{H}}\,\cos{\psi}_{{}_{H}})$ is the theoretical magnitude of the normalisation parameter.
+In the next section, we will see how the velocity of virtual particles relates the value of $A_{t}$ and the metric properties of spacetime.
+IV Spacetime metrics in the interaction domain
 To classify (pseudo-)Euclidean spaces, the so-called space index $k$ (or the index of inertia) is introduced.
@@
 The color of the dots corresponds to different values of the inelasticity $y$ as in Fig (a).
-It follows from the condition, $\beta^{*}(Q^{2}\rightarrow 0)=$1 that the normalization parameter $A_{e}=1/\sqrt{3}$.
+It follows from the condition, $\beta^{*}(Q^{2}\rightarrow 0)=$1 that the normalisation parameter $A_{e}=1/\sqrt{3}$.
 (c) The same as in panel (b), but for $\beta^{*}$ as a function of two variables $x_{{}_{Bj}}$ and $Q^{2}$ at different $y$-intervals.
@@
 Let us first discuss possible values of the parameter $A_{t}$ in formula (9), and denote by $A_{e}$ its value found from the data.
-The Standard Model assumes that the geometry of space-time known from macroscopic physics also holds in the microcosm too.
+The Standard Model assumes that the geometry of spacetime known from macroscopic physics also holds in the microcosm too.
 The derivation of the inequality (4) was based on this assumption, using the dot product of vectors in proper Euclidean space.
@@
 In this case, $A_{t}=3/\cos{\psi}_{{}_{H}}\geq 3$.
-Consequently, if it follows from an experimental data that $A_{e}\geq 3$, then the interaction of particles takes place in the region of space-time with the Minkowski geometry and the space index of $k=3$.
-The case is quite different if $A_{e}<3$. Then our assumption about the metric of space-time in the interaction region is not correct.
+Consequently, if it follows from an experimental data that $A_{e}\geq 3$, then the interaction of particles takes place in the domain of spacetime with the Minkowski geometry and the space index of $k=3$.
+The case is quite different if $A_{e}<3$. Then our assumption about the metric of spacetime in the interaction domain is not correct.
 Let us now apply the formula (9) to evaluate the speed of virtual photons.
@@
 As input, we use the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider [10].
-The magnitude of the photon virtuality, $Q^{2}$, varies over a very wide range of values.
-Neutral current interaction cross sections at low $Q^{2}\leq$100 GeV2 are dominated by virtual photon exchange.
+The magnitude of the exchange particles virtuality, $Q^{2}$, varies over a very wide range of values.
+Neutral current interaction cross sections at low $Q^{2}\leq$100 GeV2 are dominated by the virtual photon exchange.
 In the limit $Q^{2}\rightarrow 0$ GeV2 (a real photon limit), $\beta^{*}=\|\mathrm{U}_{lb}^{*}/c\|\rightarrow 1$ should hold.
@@
 The condition $\beta^{*}(Q^{2}\rightarrow 0)=$1 fixes $A_{e}$.
-In this way we get $\sqrt{3}A_{e}=1$.
+In this way one get $\sqrt{3}A_{e}=1$.
 Figures 1(c) and 1(d) show the dependence of $\beta^{*}$ on two variables, $x_{{}_{Bj}}$ and $Q^{2}$.
@@
 1 show that the speed of virtual photons at $Q^{2}>0$ GeV2 exceeds the speed of light $c$ in free space, $\beta^{*}>1$.
-Superluminal virtual photons can be interpreted as representatives of a class of superluminal particles, the tachyons [24, 25, 26, 27, 28, 29], since their properties are largely similar as shown in Table 1,
-Table 1: Comparison of the properties of tachions and virtual photons.
-Tachyon Virtual photon, $\gamma^{*}$
-mass $(m^{*})^{2}<0$ $(m^{*})^{2}=-Q^{2}<0$
-energy $\epsilon<0,>0$ $q_{0}<0,>0$
+A perusal of the litrature on faster-than-light particles reveals that virtual photons can be interpreted as representatives of a class of superluminal particles, hypothetical tachyons [26, 27, 28, 29, 30, 31], since their properties are largely similar as shown in Table 1,
+Table 1: Comparison of the properties of virtual photons and tachyons.
+Virtual photon, $\gamma^{*}$ Tachyon
+mass $(m^{*})^{2}=-Q^{2}<0$ $(m^{*})^{2}<0$
+energy $q_{0}<0,>0$ $\epsilon<0,>0$
 speed $\beta^{*}\geq 1$ $\beta^{*}\geq 1$
-$\epsilon\rightarrow 0,\beta^{*}\rightarrow\infty$ $q_{0}\rightarrow 0,\beta^{*}\rightarrow\infty$ (?)
+$q_{0}\rightarrow 0,\beta^{*}\rightarrow\infty$ (?) $\epsilon\rightarrow 0,\beta^{*}\rightarrow\infty$
+Moreover, such a result has been expected for a long time.
+Thanks to many years of extensive theoretical studies, mainly carried out by a group of Italian physicists led by E.
+Recami (for an overview see [32, 33]), it became clear that tachyons (or spacelike states) must exist as intermediate states or exchanged objects in elementary processes [34, 35, 32].
+For reviews of searches for evidence of tachyons see [36, 37].
+Thus, using a different theoretical approach and based on modern experimental data from high-energy physics, we have confirmed the predictions made more than 40 years ago.
 From the above discussion, the value of $A_{e}=1/\sqrt{3}<3$ obtained from the experimental data, contradicts the theoretical expectation, $A_{t}=3/\cos{\psi}_{{}_{H}}>3$.
@@
 In this case, the speed of the signals can be less or more than the speed of light $c$ in void and the space index change from $k=3$ to $k=2$ or even to $k=1$ and hence, $A_{t}<3$.
-This corresponds to the space-time domain, in which there are two or three time-like dimensions.
+This corresponds to the spacetime domain, in which there are two or three time-like dimensions.
 However, in this scenario there are problems not only with causality (paradoxes), the violation of unitarity and occurrence of ghost modes with negative norm (for more discussion see Refs.
@@
 II) Stochastic time.
-One can increase the pseudo-Euclideanity of space-time by adding an additional time-like dimension in the interaction region, i.e., going from signature (1,3) to signature (2,3).
-Analysis of this scenario shows [33], that in the case of a space-time with a thermally excited second time dimension the dynamics of the physical system is diffusive, not ballistic, so that all trajectories involving motion in the second time dimension are dynamically unstable, thus allowing us to avoid the difficulties outlined for nonlinear field theories.
+One can increase the pseudo-Euclideanity of spacetime by adding an additional time-like dimension in the interaction domain, i.e., going from signature (1,3) to signature (2,3).
+Analysis of this scenario shows [41], that in the case of a spacetime with a thermally excited second time dimension the dynamics of the physical system is diffusive, not ballistic, so that all trajectories involving motion in the second time dimension are dynamically unstable, thus allowing us to avoid the difficulties outlined for nonlinear field theories.
 The thermal extra time dimension behaves like an extra spatial dimension and the high temperature limit for the thermal extra time dimension is equivalent to a small compactification radius for an extra spatial dimension.
@@
 I.
-Bars with co-authors [38], [39] [32] discovered a fundamental role of the local symplectic Sp(2, R) gauge symmetry in phase space, which gave rise to new field theories in (2,4) space-time with one extra time and one extra space dimensions (2T-physics).
+Bars with co-authors [46], [47], [40] discovered a fundamental role of the local symplectic Sp(2, R) gauge symmetry in phase space, which gave rise to new field theories in (2,4) spacetime with one extra time and one extra space dimensions (2T-physics).
 This includes 2T field theories that yield one-time (1T) field theories for the Standard Model [48], General Relativity [49], SUSY [50] and others.
@@
 In the following, we give a qualitative argument for the fact that the condition $A_{t}<3$ is feasible in the framework of 2T-physics.
-First, we postulate that the interaction region of DIS has a small spatial and temporal extent, and within the framework of the 2T-physics, the interactions take place in $\mathrm{(1+1^{\prime},3+1^{\prime})}$ space with one extra time and one extra space dimensions and the gauge invariant sector of 2T-physics, namely the ghost free physical sector, effectively becomes a 1T theory with an effective 1+3 dimensions [43].
-Mathematically, the speed of a photon in the physical sector is a combination of the three effective spatial dimensions and one effective temporal dimension (the ”effective Minkowski space-time”).
-So, we assume that during the DIS process a virtual photon “captures” (or “admixes”) “new” additional spatial $x$ or time-like $\tau$ dimensions with the probability $\alpha$ and captures the “old” space-time dimensions, $\{t,x^{i}\}$, with the probability $\omega$.
+First, we postulate that the interaction domain of DIS has a small spatial and temporal extent, and within the framework of the 2T-physics, the interactions take place in $\mathrm{(1+1^{\prime},3+1^{\prime})}$ space with one extra time and one extra space dimensions and the gauge invariant sector of 2T-physics, namely the ghost free physical sector, effectively becomes a 1T theory with an effective 1+3 dimensions [51].
+Mathematically, the speed of a photon in the physical sector is a combination of the three effective spatial dimensions and one effective temporal dimension (the ”effective Minkowski spacetime”).
+So, we assume that during the DIS process a virtual photon “captures” (or “perceives”) the “new” additional spatial $x$ or time-like $\tau$ dimensions with the probability $\alpha$ and captures the “old” spacetime dimensions, $\{t,x^{i}\}$, with the probability $\omega$.
 We label the components containing the conventional time dimension with index $H$, and the components containing an additional time-like dimension with index $B$.
@@
 Here, $A_{H}=3/\cos{\psi}_{{}_{H}}$ and $A_{B}=3/\cos{\psi}_{{}_{B}}$.
-In the limit $Q^{2}\to 0$, all speeds in Eq. (15) should be equated to unity. As a result, we get the normalization condition $\sqrt{3}A_{eff}=1$,
+In the limit $Q^{2}\to 0$, all speeds in Eq. (15) should be equated to unity. As a result, we get the normalisation condition $\sqrt{3}A_{eff}=1$,
 $1=3\sqrt{3}\Big{[}\frac{1-\alpha-3\alpha^{2}}{\cos{\psi}_{{}_{H}}}+\frac{\alpha+3\alpha^{2}}{\cos{\psi}_{{}_{B}}}\Big{]}.$ (16)
@@
 These restrictions on $\cos{\psi}_{{}_{H}}$ and $\cos{\psi}_{{}_{B}}$ are quite soft.
-Thus, in the space-time volume of a high-energy reaction a “mixing in” of extra dimensions is possible and the normalization condition $\sqrt{3}A_{eff}=1$ is feasible.
+Thus, in the spacetime volume of a high-energy reaction a “mixing in” of extra dimensions is possible and the normalisation condition $\sqrt{3}A_{eff}=1$ is feasible.
 The specific values of $\cos{\psi}_{{}_{H}}$, $\cos{\psi}_{{}_{B}}$, $\alpha$, ${\beta^{2}}_{H}$, $\tilde{\beta^{2}}_{H}$, $\tilde{u^{2}}_{B}$, $\tilde{u^{2}}_{B}$ can be found by fitting Eq.
@@
 V Conclusions
-In this Letter, we have developed the method based on the quantum uncertainty relations and the theory of indirect measurements, which allows to evaluate the module of the group velocity of virtual photons $\gamma^{*}$ and to reveal the type of space-time geometry in the interaction region of high-energy deep inelastic scattering processes.
-The HERA data show that the speed of virtual photons, $\beta^{*}$, exceeds the speed of light in free space, $\beta^{*}\geq 1$.
+This communication presents formula (9) for estimating the lower bound of the modulus of the group velocity of a virtual particle.
+When analysing experimental data, the uncertainties included in the formula should be calculated by methods of indirect measurement theory.
+To estimate the speed of virtual photons, the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider were used as input.
+The HERA data at $Q^{2}\leq$100 GeV2 show that the normalized speed of virtual photons, $\beta^{*}$, exceeds the speed of light in free space, $\beta^{*}\geq 1$.
 The superluminal speed means that the virtual photons $\gamma^{*}$ behave tachyon-like.
-A solution of the problem of the normalization condition $\beta^{*}=1$ at $Q^{2}=0\,\rm{GeV}^{2}$ is proposed in the framework of ideas of ”2T-physics” developed by I.
+This characteristic behavior of virtual photons is consistent with the predictions made by Recami and co-authors in the 1970s and 1980s that tachyons are intermediate states or exchanged objects in elementary processes.
+When the normalisation condition $\beta^{*}=1$ is imposed in a real photon limit, the deep relationship between the normalisation parameter $A_{t}$ and the type of spacetime geometry in the interaction domain is revealed.
+A solution of the problem of the normalisation condition $\beta^{*}=1$ at $Q^{2}=0\,\rm{GeV}^{2}$ is proposed in the framework of ideas of ”Two-time physics” developed by I.
 Bars.
-It is shown that the “admixture” of extra time and extra space dimensions in DIS processes makes it possible to satisfy the normalization condition $\sqrt{3}A_{eff}=1$ as observed in the effective physical (1,3) Minkowski space-time.
-The experimental results for virtual photons presented here are based on only a small fraction of the HERA data to illustrate the method, but the analysis can be extended to the full data set, and these investigations are left to future work.
+It is shown that by admixing one extra time and one exta space dimensions to 4D spacetime in the domain of DIS processes, it is possible to satisfy the normalisation condition $\sqrt{3}A_{eff}=1$ as observed in the effective physical (1,3) Minkowski spacetime.
+For the sake of illustration of the method, the results for virtual photons presented here are based on only a small fraction of the HERA data, but the analysis can be extended to the full data set, and these investigations are left to future work.
 ACKNOWLEDGMENTS
@@
 Nikolaev and B.S.
-Yuldashev for their mentorship in the early days.
+Yuldashev for their mentorship in the early days of this amazing scientific journey.
 The author expresses his gratitude to the Ministry of Science and Education of the Russian Federation and the DESY Directorate for their long-term financial support and for the hospitality extended to the non-DESY members of the ZEUS collaboration.